9x - (5x-5) = 37
How do I solve and what is the x

Answer by mananth(16946)

9x - (5x-5) = 37
First step is to remove the parenthesis
whenever there is negative sign before the parenthesis all the signs of the terms in the parenthesis become opposite.
9x-5x+5 = 37
4x+5 = 37
-5 -5
4x= 37-5
4x=32
/4
x= 8
...
m.ananth@hotmail.ca